OpenStudy (anonymous):

Factor the following polynomial: 1. 2a+2b+wa+wb 2. pr-2r-ap+2a 3. -3a+3+ax-x 4. 4y^2-49 5. x^2y^2-9c^2 6. x^2-20x+100 7. 9t^2+30t+25 8. x^3-64 9. 27w^3-125 10. x^2y-2x^2-y+2

OpenStudy (akashdeepdeb):

Use the following properties. \[cp + cq + dp + dq = c(p + q) + d(p+q) = (c+d)(p+q)\] \[a^2-b^2 = (a+b)(a-b)\] \[a^2 - 2ab + b^2 = {(a-b)}^2\] \[a^2 + 2ab + b^2 = {(a+b)}^2\] \[x^3-y^3 = (x-y)(x^2+xy+y^2)\] \[x^3 + y^3 = (x+y)(x^2-xy+y^2)\]

OpenStudy (triciaal):

1. 2(a + b ) + w(a + b) = (2 + w )(a + b) use the properties given by @AkashdeepDeb

OpenStudy (lxelle):

In Openstudy, we don't give answers, so don't expect us to do so. We are her to guide you with your questions. You need to do revision first before doing those exrcises/assignments.